OpenStudy (anonymous):

log2x-log2(x-2)=3 find x

OpenStudy (mathstudent55):

Use this rule of logs first. \( \large \log_b x - \log_b y = \log_b \dfrac{x}{y} \)

OpenStudy (whpalmer4):

@666slayer are you stuck here?

OpenStudy (anonymous):

yes the answer to x is 16/7 but i don't know how to work the problem to get the answer.

OpenStudy (whpalmer4):

Okay, doing as my colleague suggested: \[\log_2 x - \log_2 (x-2) = 3\]\[\log_2 (\frac{x}{x-2}) = 3\]

OpenStudy (whpalmer4):

Do you see how to do it now?

OpenStudy (whpalmer4):

Raise the base of the logarithm to both sides…\[2^{\log_2 (\frac{x}{x-2})} = 2^3\]\[b^{\log_b u} = u\]

OpenStudy (anonymous):

2log2(x/x-2)=8 where do you go from here?

OpenStudy (whpalmer4):

\[2^{\log_2 (\frac{x}{x-2})} = 2^3\]But we apply that principle \[b^{\log_b u} = u\]and we get \[\frac{x}{x-2} = 8\]You can solve that, right?

OpenStudy (whpalmer4):

\[x = 8(x-2)\]\[x=8x-16\]\[-7x=-16\]\[x=\frac{16}{7}\]

OpenStudy (anonymous):

Thanks a lot i got it now